the distance moved by a body at the end of t1=3s and t2=5s are 126m and 250m respectively. what is the initial velocity of the body.

1
Expert's answer
2022-01-24T01:13:28-0500

"s(t) = v_0t +\frac{at^2}{2}"

"s(3)= 3v_0+4.5a =126(1)"

"s(5)= 5v_0+12.5a = 250(2)"

"\text{From (1) and (2)}"

"v_0=27\frac{m}{s}"


"\text{Answer:} 27\frac{m}{s}"